# Finite Neural Networks as Mixtures of Gaussian Processes: From Provable Error Bounds to Prior Selection

### v1

@@ 3.2 Wasserstein Distance
[…]
${\mathbb{W}}_{\rho}(p,q)\coloneqq\left(\inf_{\gamma\in\Gamma(p,q)}{\mathbb{E}}_{({{x}},{{y}})\sim\gamma}[\|{{x}}-{{y}}\|^{\rho}]\right)^{\frac{1}{\rho}}=\left(\inf_{\gamma\in\Gamma(p,q)}\int_{{\mathbb{R}}^{d}\times{\mathbb{R}}^{d}}\|{{x}}-{{y}}\|^{\rho}\gamma(d{{x}},d{{y}})\right)^{\frac{1}{\rho}},$ (1)
@@
where $\Gamma(p,q)\subset{\mathcal{P}}_{\rho}({\mathbb{R}}^{d}\times{\mathbb{R}}^{d})$ represents the set of probability distributions with marginal distributions $p$ and $q$. It can be shown that ${\mathbb{W}}_{\rho}$ is a metric, which is given by the minimum cost, according to the $\rho-$power of the Euclidean norm, required to transform one probability distribution into another (Villani et al., 2009). Furthermore, another attractive property of the $\rho$-Wasserstein distance, which distinguishes it from other divergence measures such as the KL divergence (Hershey and Olsen, 2007), is that closeness in the $\rho$-Wasserstein distance implies closeness in the first $\rho$ moments. This result is formalized in Lemma 2 below
Lemma 2
For $p,q\in{\mathcal{P}}_{2}({\mathbb{R}}^{n})$ and $\rho\in\mathbb{N}$ it holds that
$|{\mathbb{E}}_{{{x}}\sim p}[\|{{x}}\|^{\rho}]^{\frac{1}{\rho}}-{\mathbb{E}}_{{{y}}\sim q}[\|{{y}}\|^{\rho}]^{\frac{1}{\rho}}|\leq{\mathbb{W}}_{\rho}(p,q)$ (2)
Proof  Let $\delta_{0}$ be the Dirac measure centered at zero, then, by the triangle inequality, it holds that ${\mathbb{W}}_{\rho}(p,\delta_{0})\leq{\mathbb{W}}_{\rho}(p,q)+{\mathbb{W}}_{\rho}(q,\delta_{0})$ and ${\mathbb{W}}_{\rho}(q,\delta_{0})\leq{\mathbb{W}}_{\rho}(q,p)+{\mathbb{W}}_{\rho}(p,\delta_{0})$. Additionally, using the symmetry axiom of a distance metric, we have that ${\mathbb{W}}_{\rho}(q,p)={\mathbb{W}}_{\rho}(p,q)$. Then, these inequalities can be combined to $|{\mathbb{W}}_{\rho}(p,\delta_{0})-{\mathbb{W}}_{\rho}(q,\delta_{0})|\leq{\mathbb{W}}_{\rho}(p,q)$. The proof is concluded by noticing that, as shown in Villani et al. (2009), it holds that that ${\mathbb{W}}_{\rho}(p,\delta_{0})={\mathbb{E}}_{{{x}}\sim p}[\|{{x}}\|^{\rho}]^{\frac{1}{\rho}}$ and ${\mathbb{W}}_{\rho}(q,\delta_{0})={\mathbb{E}}_{{{x}}\sim q}[\|{{x}}\|^{\rho}]^{\frac{1}{\rho}}$.
@@
As is common in the literature, in what follows, we will focus on the 2-Wasserstein distance. However, it is important to note that since ${\mathbb{W}}_{1}\leq{\mathbb{W}}_{2}$, the methods presented in this work naturally extend to 1-Wasserstein distance. Detailed comparisons and potential improvements when utilizing the $1$-Wasserstein distance are reported in the Appendix.
@@ 4.1 Stochastic Neural Networks (SNNs)
For an input vector $x\in{\mathbb{R}}^{n_{0}}$, we consider a fully connected neural network of $K$ hidden layers $f^{w}(x)$ defined iteratively as follows for $k\in\{0,\ldots,K\}$:
${{{z}}}_{0}={{x}},$ $L_{w_{k}}({{{z}}}_{k})={{{W}}}_{k}{{{z}}}_{k}+{{{b}}}_{k},$ ${{{z}}}_{k+1}=\begin{cases}L_{{{w}}_{k}}({{{z}}}_{k})&\text{if}\ k=0,\\ L_{{{w}}_{k}}({\sigma}_{k}({{{z}}}_{k}))&\text{otherwise,}\end{cases}$ $f^{{{w}}}(x)={{{z}}}_{K+1},$ (3)
@@
where, for $n_{k}$ being the number of neurons of layer $k$, we have that ${\sigma}_{k}:{\mathbb{R}}^{n_{k}}\to{\mathbb{R}}^{n_{k}}$ is the vector of piecewise continuous activation functions (one for each neuron), ${{{W}}}_{k}\in{\mathbb{R}}^{n_{k+1}\times n_{k}}$ and ${{{b}}}_{k}\in{\mathbb{R}}^{n_{k+1}}$ are the matrix of weights and vector of bias parameters of the $k-$th layer of the network. We denote by ${{{w}}}_{k}=\{{{{W}}}_{k},{{{b}}}_{k}\}$ the union of all parameters of the $k$-th layer and by ${{{w}}}=\{{{{w}}}_{k}\}_{k=0}^{K}$ the union of all the neural network parameters, which we simply call weights or parameters of the neural network. $L_{{{w}}_{K}}({{{z}}}_{K})$ is the final output of the network, possibly corresponding to the vector of logits in case of classification problems.
@@
In this work, we assume that the weights ${{{w}}}$, rather than being fixed, are distributed according to a probability distribution $p_{{{{w}}}}$. [^3] This includes architectures with both deterministic and stochastic weights, where the distribution over deterministic weights can be modeled as a Dirac delta function. For any $x\in\mathbb{R}^{n_{0}},$ placing a distribution over the weights leads to a distribution over the outputs of the NN. That is, $f^{w}(x)$ is a random variable and $\{f^{w}(x)\}_{x\in\mathbb{R}^{n_{0}}}$ is a stochastic process, which we call a stochastic neural network (SNN) to emphasize its randomness (Yu et al., 2021). In particular, $f^{w}(x)$ follows a probability distribution $p_{nn(x)}$, which, as shown in Adams et al. (2023), can be iteratively defined over the layers of the SNN. Specifically, as shown in Eqn. (4) below, $p_{nn(x)}$ is obtained by recursively propagating through the layers of the NN architecture the distribution induced by the random weights at each layer. The propagation is obtained by marginalization of the output distribution at each layer with the distribution of the previous layer.
$p_{nn({{x}}),1}={\mathbb{E}}_{{{{z}}}_{0}\sim\delta_{{x}}}[p_{L_{{{w}}_{0}}({{{z}}}_{0})}]$
$p_{nn({{x}}),k+1}={\mathbb{E}}_{{{z}}_{k}\sim p_{nn({{x}}),k}}[p_{L_{{{w}}_{k}}({\sigma}({{z}}_{k}))}],\qquad k\in\{1,...,K\}$ (4)
$p_{nn(x)}=p_{nn({{x}}),K+1},$
@@
where $p_{L_{{{w}}_{k}}({{{z}}})}=L_{{{w}}_{k}}({{{z}}})\#p_{{{w}}_{k}}$ and $\delta_{{x}}$ is the Delta Dirac function centered at $x$. For any finite subset ${\mathcal{X}}\subset{\mathbb{R}}^{n_{0}}$ of input points, we use $p_{nn({\mathcal{X}})}$ to denote the joint distribution of the output of $f^{w}$ evaluated at the points in ${\mathcal{X}}$.
@@
In what follows, because of its practical importance and the availability of closed-form solutions, we will introduce our methodological framework under the assumption that $p_{{{{w}}}}$ is a multivariate Gaussian distribution that is layer- or neuron-wise correlated, i.e., $p_{{w}}=\mathcal{N}({m}_{{w}},\Sigma_{{w}})$ with $\Sigma_{{w}}=\mathrm{diag}(\Sigma_{{{w}}_{0}},\ldots,\Sigma_{{{w}}_{K}})$ and in the case of neuron-wise correlation $\Sigma_{{{w}}_{k}}=\mathrm{diag}(\Sigma_{{{w}}_{k,1}},\ldots,\Sigma_{{{w}}_{k,n_{k}}})$. We stress that this does not imply that the output distribution of the SNN, $p_{nn}(x)$, is also Gaussian; in fact, in general, it is not. Furthermore, we should also already remark that, as we will explain in Subsection 6.1 and illustrate in the experimental results, the methods we present in this paper can be extended to more general (non-necessarily Gaussian) $p_{w}\in{\mathcal{P}}_{2}(\cdot)$.
@@
The above definition of SNNs encompasses several stochastic models of importance for applications, including Bayesian Neural Networks (BNNs) trained with Gaussian Variational Inference methods such as, e.g., Bayes by Backprop (Blundell et al., 2015) or Noisy Adam (Zhang et al., 2018) [^4] The methods presented in this paper can be applied to other approximate inference methods such as HMC (Neal, 2012), for which $p_{w}$ takes the form of a categorical distribution, as explained in Subsection 6.1., NNs with only a subset of stochastic layers (Tang and Salakhutdinov, 2013), Gaussian Dropout NNs (Srivastava et al., 2014), and NNs with randomized smoothing and/or stochastic inputs (Cohen et al., 2019). The methods proposed in this paper apply to all of them. In particular, in the case of BNNs, in Section 8, we will show how our approach can be used to investigate both the prior and posterior behavior, in which case, depending on the context, $p_{w}$ could represent either the prior or the posterior distribution of the weights and biases.
@@ 4.2 Problem Statement
${\mathbb{W}}_{2}(p_{nn(\mathcal{X})},q_{nn({\mathcal{X}})})\leq\epsilon.$ (5)
@@
While in Problem 1 we seek for a Guassian approximation of a SNN, one could also consider the complementary problem of finding the parameters of a SNN that best approximate a given GMM. Such a problem also has wide application. In fact, a solution to this problem would allow one to encode informative functional priors represented via Gaussian processes to SNNs, addressing one of the main challenges in performing Bayesian inference with neural networks (Flam-Shepherd et al., 2017; Tran et al., 2022). As the ${\mathbb{W}}_{2}$ satisfies the triangle inequality and because there exist closed-form expressions for an upper-bound on the Wasserstein distance between Gaussian Mixture distributions (Delon and Desolneux, 2020), a solution to Problem 1 can be readily used to address the above mentioned complementary problem. This will be detailed in Section 7.2 and empirically demonstrated in Section 8.3.
@@
Note that in Problem 1 we consider a GMM of arbitrary size. This is because such a model can approximate any continuous distribution arbitrarily well, and, consequently, guarantees of convergence to satisfy Problem 1 can be obtained, as we will prove in Section 6. However, Problem 1 could be restricted to the single Gaussian case, in which case Problem 1 reduces to finding the GP that best approximates a SNN. However, in this case, because $p_{nn(\mathcal{X})}$ is not necessarily Gaussian, the resulting distance bound between the GP and the SNN may not always be made arbitrarily small.
@@ 4.2.1 Approach Outline
Our approach to solving Problem 1 is based on iteratively approximating the output distribution of each layer of a SNN as a mixture of Gaussian distributions and quantifying the approximation error introduced by each operation. As illustrated in Figure 2, following the definition of $p_{nn}({{x}})$ in Eqn (4), the first step is to perform a linear combination of the input point ${{x}}$ with the parameters of the first layer (step a). Because the SNN weights are assumed to be jointly Gaussian and jointly Gaussian random variables are closed under linear combination, this linear operation leads to a Gaussian distribution. Then, before propagating this distribution through an activation function, a signature operation on the resulting distribution is performed; that is, the continuous distribution is approximated into a discrete distribution (step b). After the signature approximation operation, the resulting discrete distribution is passed through the activation function (step c) and a linear combination with the weights of the next layer is performed (step d). Under the assumption that the weights are jointly Gaussian, this linear operation results into a Gaussian mixture distribution of size equal to the support of the discrete distribution resulting from the signature operation. To limit the computational burden of propagating a Gaussian mixture with a large number of components, a compression operation is performed that compresses this distribution into another mixture of Gaussian distributions with at most $M$ components for a given $M$ (step e). After this, a signature operation is performed again on the resulting Gaussian mixture distribution (step f), and the process is repeated until the last layer. Consequently, to construct $q_{nn({{x}})}$, the Gaussian mixture approximation of $p_{nn({{x}})}$, our approach iteratively performs four operations: linear transformation, compression, signature approximation, and propagation through an activation function. To quantify the error in our approach, we derive formal error bounds in terms of the Wasserstein distance for each of these operations and show how the resulting bounds can be combined via interval arithmetic to an upper bound of ${\mathbb{W}}_{2}(p_{nn({{x}})},q_{nn({{x}})})$.

### v2

@@ 3.2 Wasserstein Distance
To approximate SNNs to GMMs and vice-versa, and quantify the quality of the approximation, we need a notion of distance between probability distributions. While various distance measures are available in the literature, in this work we consider the Wasserstein distance (Gibbs and Su, 2002). To define the Wasserstein distance, for $\rho\geq 1$ we define the $\rho$-Wasserstein space of distributions ${\mathcal{P}}_{\rho}({\mathbb{R}}^{d})$ as the set of probability distributions with finite moments of order $\rho$, i.e., any $p\in{\mathcal{P}}_{\rho}({\mathbb{R}}^{d})$ is such that $\int_{{\mathbb{R}}^{d}}\|x\|^{\rho}dp(x)<\infty$. For $p,q\in{\mathcal{P}}_{\rho}({\mathbb{R}}^{d})$, the $\rho$-Wasserstein distance ${\mathbb{W}}_{\rho}$ between $p$ and $q$ is defined as
${\mathbb{W}}_{\rho}(p,q)\coloneqq\left(\inf_{\gamma\in\Gamma(p,q)}{\mathbb{E}}_{({{x}},{{z}})\sim\gamma}[\|{{x}}-{{z}}\|^{\rho}]\right)^{\frac{1}{\rho}}=\left(\inf_{\gamma\in\Gamma(p,q)}\int_{{\mathbb{R}}^{d}\times{\mathbb{R}}^{d}}\|{{x}}-{{z}}\|^{\rho}\gamma(d{{x}},d{{z}})\right)^{\frac{1}{\rho}},$ (1)
@@
where $\Gamma(p,q)\subset{\mathcal{P}}_{\rho}({\mathbb{R}}^{d}\times{\mathbb{R}}^{d})$ represents the set of probability distributions with marginal distributions $p$ and $q$. It can be shown that ${\mathbb{W}}_{\rho}$ is a metric, which is given by the minimum cost, according to the $\rho$-power of the Euclidean norm, required to transform one probability distribution into another (Villani et al., 2009). Furthermore, another attractive property of the $\rho$-Wasserstein distance, which distinguishes it from other divergence measures such as the KL divergence (Hershey and Olsen, 2007), is that closeness in the $\rho$-Wasserstein distance implies closeness in the first $\rho$ moments. [^3] See Lemma 28 in Appendix A.
@@
In what follows, we focus on the $2$-Wasserstein distance. [^4] Since ${\mathbb{W}}_{1}\leq{\mathbb{W}}_{2}$, the methods presented in this work naturally extend to the 1-Wasserstein distance. Detailed comparisons and potential improvements when utilizing the $1$-Wasserstein distance are reported in the Appendix. While closed-form expressions exist for the $2$-Wasserstein distance for Gaussian distributions (Givens and Shortt, 1984),555 For $p_{1}=\mathcal{N}({m}_{1},\Sigma_{1})\in\mathcal{G}({\mathbb{R}}^{n})$ and $p_{2}=\mathcal{N}({m}_{2},\Sigma_{2})\in\mathcal{G}({\mathbb{R}}^{n})$, it holds that ${\mathbb{W}}_{2}^{2}(p_{1},p_{2})=\|{m}_{1}-{m}_{2}\|^{2}+\mathrm{trace}\left(\Sigma_{1}+\Sigma_{2}-2\left(\Sigma_{1}^{1/2}\Sigma_{2}\Sigma_{1}^{1/2}\right)^{1/2}\right)$. this is not the case for Gaussian mixture distributions. However, a tractable upper bound for the ${\mathbb{W}}_{2}$ is provided by the ${\mathbb{MW}}_{2}$ distance, formally defined below following (Delon and Desolneux, 2020).
Definition 2 (${\mathbb{MW}}_{2}$ Distance)
Let $p=\sum_{i=1}^{N_{p}}{{{\pi}}_{p}^{{(i)}}}p_{i}\in\mathcal{G}_{N_{1}}({\mathbb{R}}^{n})$ and $q=\sum_{j=1}^{N_{q}}{{{\pi}}_{q}^{{(j)}}}q_{j}\in\mathcal{G}_{N_{1}}({\mathbb{R}}^{n})$ be two Gaussian Mixture distributions. Then, the ${\mathbb{MW}}_{2}$-distance is defined as
${\mathbb{MW}}_{2}(p,q)\coloneqq\left(\inf_{\gamma\in\Gamma(p,q)\cap\mathcal{G}_{<\infty}({\mathbb{R}}^{n}\times{\mathbb{R}}^{n})}{\mathbb{E}}_{({{x}},{{z}})\sim\gamma}[\|{{x}}-{{z}}\|^{2}]\right)^{\frac{1}{2}}=\bigg{(}\min_{{{\pi}}\in\Gamma({{\pi}}_{p},{{\pi}}_{q})}\sum_{i,j}{{{\pi}}^{{(i,j)}}}{\mathbb{W}}_{2}^{2}(p_{i},q_{j})\bigg{)}^{\frac{1}{2}}$ (2)
where $\mathcal{G}_{<\infty}({\mathbb{R}}^{n}\times{\mathbb{R}}^{n})$ is the set of Gaussian mixture distributions with finitely many components.
The ${\mathbb{MW}}_{2}$ distance is a Wasserstein-type distance that restricts the set of possible couplings to Gaussian mixtures. This guarantees that ${\mathbb{MW}}_{2}$ computation reduces to a finite discrete linear program with $N_{p}\cdot N_{q}$ optimization variables and coefficients (Delon and Desolneux, 2020). The coefficients are given by the Wasserstein distance between the mixture’s Gaussian components, which have closed-form expressions. Furthermore, as Definitions 1 and 2 only differ for the intersection on $\mathcal{G}_{<\infty}$ in the couplings, it is easy to see that ${\mathbb{MW}}_{2}(p,q)\geq\mathbb{W}_{2}(p,q)$.
@@ 4.1 Stochastic Neural Networks (SNNs)
For an input vector $x\in{\mathbb{R}}^{n_{0}}$, we consider a feedforward neural network of $K$ hidden layers $f^{w}(x)$ defined iteratively as:
${{z}}_{1}=L^{{{w}}_{0}}({{x}})$ (3a)
${{z}}_{k+1}=L^{{{w}}_{k}}({\sigma}({{z}}_{k})),\qquad k\in\{1,\ldots,K\}$ (3b)
$f^{{{w}}}({{x}})={{{z}}}_{K+1}$ (3c)
@@
where, for $n_{k}$ being the number of neurons of layer $k$, we have that ${\sigma}_{k}:{\mathbb{R}}^{n_{k}}\to{\mathbb{R}}^{n_{k}}$ is a the vector of piecewise continuous activation functions, and $L^{{{w}}_{k}}:{\mathbb{R}}^{n_{k}}\rightarrow{\mathbb{R}}^{n_{k+1}}$ denotes an affine transformation, such as a fully-connected (dense) layer or a convolutional layer. Each $L^{{{w}}_{k}}$ is parametrized by weights ${{W}}_{k}$, e.g., a transformation matrix or convolutional kernel, and bias ${{b}}_{k}$, which are collectively stored in ${{w}}_{k}=\{{{W}}_{k},{{b}}_{k}\}$. We denote by ${{w}}=\{{{w}}_{k}\}_{k=0}^{K}$ the union of all the neural network parameters. ${{z}}_{K+1}$} is the final output of the network, possibly corresponding to the vector of logits in case of classification problems.
@@
In this work, we assume that the parameters ${{w}}_{k}$ of each layer $k$, rather than being fixed, are distributed according to a probability distribution $p_{{{w}}_{k}}$. [^6] This includes architectures with both deterministic and stochastic parameters, where the distribution over deterministic parameters can be modeled as a Dirac delta function. For any $x\in\mathbb{R}^{n_{0}},$ placing a distribution over the parameters leads to a distribution over the outputs of the NN. That is, $f^{w}(x)$ is a random variable and $\{f^{w}(x)\}_{x\in\mathbb{R}^{n_{0}}}$ is a stochastic process, which we call a stochastic neural network (SNN) to emphasize its randomness (Yu et al., 2021). In particular, $f^{w}(x)$ follows a probability distribution $p_{nn(x)}$, which, as shown in Adams et al. (2023), can be iteratively defined over the layers of the SNN. Specifically, as shown in Eqn (4) below, $p_{nn(x)}$ is obtained by recursively propagating through the layers of the NN architecture the distribution induced by the random parameters at each layer. The propagation is obtained by marginalization of the output distribution at each layer with the distribution of the previous layer:
$p_{nn({{x}}),1}={\mathbb{E}}_{{{{z}}}_{0}\sim\delta_{{x}}}[L^{{{w}}_{0}}({{{z}}}_{0})\#p_{{{w}}_{0}}]$ (4a)
$p_{nn({{x}}),k+1}={\mathbb{E}}_{{{z}}_{k}\sim p_{nn({{x}}),k}}[L^{{{w}}_{k}}({\sigma}({{z}}_{k}))\#p_{{{w}}_{k}}],\qquad k\in\{1,\ldots,K\}$ (4b)
$p_{nn(x)}=p_{nn({{x}}),K+1}$ (4c)
@@
where $\delta_{{x}}$ is the Delta Dirac function centered at xx. [^7] Equivalently, for all $k$, $p_{nn({{x}}),k+1}$ is a mixture distribution with infinitely many components of the form $L^{{{w}}_{k}}(\sigma({{z}}_{k}))\#p_{{{w}}_{k}}$ for ${{z}}_{k}\in{\mathbb{R}}^{n_{k}}$, weighted by $p_{nn({{x}}),k}$, or, equivalently, the pushforward of the joint distribution of $p_{{{w}}_{k}}$ and $p_{nn({{x}}),k}$ by $L^{{{w}}_{k}}(\sigma({{z}}_{k}))$. For any finite subset ${\mathcal{X}}\subset{\mathbb{R}}^{n_{0}}$ of input points, we use $p_{nn({\mathcal{X}})}$ to denote the joint distribution of the output of $f^{w}$ evaluated at the points in $\mathcal{X}$}.
@@
In what follows, because of its practical importance and the availability of closed-form solutions, we will introduce our methodological framework under the assumption that $p_{{{w}}_{k}}$ is a multivariate Gaussian distributions for all $k$. We stress that this does not imply that the output distribution of the SNN, $p_{nn}(x)$, is also Gaussian; in fact, in general, it is not. Furthermore, we should also already remark that, as we will explain in Remark 16 and illustrate in the experimental results, the methods we present in this paper can be extended to more general (non-necessarily Gaussian) ${p_{{{w}}_{k}}}\in{\mathcal{P}}_{2}(\cdot)$.
@@
The above definition of SNNs encompasses several stochastic models of practical importance, including Bayesian Neural Networks (BNNs) trained with Gaussian Variational Inference methods such as, e.g., Bayes by Backprop (Blundell et al., 2015) or Noisy Adam (Zhang et al., 2018), NNs with only a subset of stochastic layers (Tang and Salakhutdinov, 2013), Gaussian Dropout NNs (Srivastava et al., 2014), and NNs with randomized smoothing and/or stochastic inputs (Cohen et al., 2019). The methods proposed in this paper apply to all such models that assume the weight distributions of different layers are independent. In particular, in the case of BNNs, in Section 8, we will show how our approach can be used to investigate both the prior and posterior behavior, in which case, depending on the context, $p_{w}$ could represent either the prior or the approximate posterior distribution of the weights and biases.
@@ 4.2 Problem Statement
${\mathbb{W}}_{2}\left(p_{nn(\mathcal{X})},q_{nn({\mathcal{X}})}\right)\leq\epsilon.$ (5)
@@
While in Problem 1 we seek for a Guassian approximation of a SNN, one could also consider the complementary problem of finding the parameters of a SNN that best approximate a given GMM. Such a problem also has wide application. In fact, a solution to this problem would allow one to encode informative functional priors represented via Gaussian processes to SNNs, addressing one of the main challenges in performing Bayesian inference with neural networks (Flam-Shepherd et al., 2017; Tran et al., 2022). As the ${\mathbb{W}}_{2}$ satisfies the triangle inequality and because there exist closed-form expressions for an upper bound on the Wasserstein distance between Gaussian Mixture distributions (Delon and Desolneux, 2020), a solution to Problem 1 can be readily used to address the above mentioned complementary problem. This will be detailed in Section 7.2 and empirically demonstrated in Section 8.3.
@@
Note that in Problem 1 we consider a GMM of arbitrary finite size. This is because such a model can approximate any continuous distribution arbitrarily well, and, consequently, guarantees of convergence to satisfy Problem 1 can be obtained, as we will prove in Section 6. However, Problem 1 could be restricted to the single Gaussian case, in which case Problem 1 reduces to finding the GP that best approximates a SNN. However, in this case, because $p_{nn(\mathcal{X})}$ is not necessarily Gaussian, the resulting distance bound between the GP and the SNN may not always be made arbitrarily small.
@@ 4.2.1 Approach Outline
Our approach to solving Problem 1 is based on iteratively approximating the output distribution of each layer of a SNN as a mixture of Gaussian distributions and quantifying the approximation error introduced by each operation. As illustrated in Figure 2, following the definition of $p_{nn({{x}})}$ in Eqn (4), the first step is to perform an affine combination of the input point x{{x}} with the parameters of the first layer (step a). Because the SNN parameters are assumed to be jointly Gaussian and jointly Gaussian random variables are closed under affine combination, this affine operation leads to a Gaussian distribution. Then, before propagating this distribution through an activation function, a signature operation on the resulting distribution is performed; that is, the continuous distribution is approximated into a discrete distribution (step b). After the signature approximation operation, the resulting discrete distribution is passed through the activation function (step c) and an affine combination with the parameters of the next layer is performed (step d). Under the assumption that the parameters are jointly Gaussian, this affine operation results into a Gaussian mixture distribution of size equal to the support of the discrete distribution resulting from the signature operation. To limit the computational burden of propagating a Gaussian mixture with a large number of components, a compression operation is performed that compresses this distribution into another mixture of Gaussian distributions with at most $M$ components for a given $M$ (step e). After this, a signature operation is performed again on the resulting Gaussian mixture distribution (step f), and the process is repeated until the last layer. Consequently, to construct $q_{nn({{x}})}$, the Gaussian mixture approximation of $p_{nn({{x}})}$, our approach iteratively performs four operations: affine transformation, compression, signature approximation, and propagation through an activation function. To quantify the error in our approach, we derive formal error bounds in terms of the Wasserstein distance for each of these operations and show how the resulting bounds can be combined via interval arithmetic to an upper bound of ${\mathbb{W}}_{2}\left(p_{nn({{x}})},q_{nn({{x}})}\right)$.